Determine the unique solution of the following differential equations by using Laplace transforms:

y''(t) + 2y'(t) -3y(t) = e-3t if y'(0) = 0 and y(0) = 0


1
Expert's answer
2022-04-14T09:22:55-0400

"y''\\left( t \\right) +2y'\\left( t \\right) -3y\\left( t \\right) =e^{-3t},y'\\left( 0 \\right) =0,y\\left( 0 \\right) =0\\\\y\\left( t \\right) \\rightarrow Y\\left( p \\right) \\\\p^2Y\\left( p \\right) -py\\left( 0 \\right) -y'\\left( 0 \\right) +2\\left( pY\\left( p \\right) -y\\left( 0 \\right) \\right) -3Y\\left( p \\right) =\\frac{1}{p+3}\\\\\\left( p^2+2p-3 \\right) Y\\left( p \\right) =\\frac{1}{p+3}\\\\Y\\left( p \\right) =\\frac{1}{\\left( p+3 \\right) ^2\\left( p-1 \\right)}\\\\\\frac{1}{\\left( p+3 \\right) ^2\\left( p-1 \\right)}=\\frac{A}{\\left( p+3 \\right) ^2}+\\frac{B}{p+3}+\\frac{C}{p-1}\\\\1=A\\left( p-1 \\right) +B\\left( p+3 \\right) \\left( p-1 \\right) +C\\left( p+3 \\right) ^2\\\\\\left( B+C \\right) p^2+\\left( A+2B+6C \\right) p+\\left( -A-3B+9C \\right) =1\\\\\\left\\{ \\begin{array}{c}\tB+C=0\\\\\tA+2B+6C=0\\\\\t-A-3B+9C=1\\\\\\end{array} \\right. \\Rightarrow A=-\\frac{1}{4},B=-\\frac{1}{16},C=\\frac{1}{16}\\\\Y\\left( p \\right) =-\\frac{1}{4\\left( p+3 \\right) ^2}-\\frac{1}{16\\left( p+3 \\right)}+\\frac{1}{16\\left( p-1 \\right)}\\mapsto -\\frac{1}{4}te^{-3t}-\\frac{1}{16}e^{-3t}+\\frac{1}{16}e^t\\\\y\\left( t \\right) =-\\frac{1}{4}te^{-3t}-\\frac{1}{16}e^{-3t}+\\frac{1}{16}e^t"